OpenStudy (frizzy.franny):

You accidentally slide a glass of milk off a table that is 0.9 m tall. How long does it take the milk to hit the ground?

OpenStudy (irishboy123):

ideas?

OpenStudy (frizzy.franny):

answer choices are A. 0.43 s B. 0.31 s C. 0.20 s D. 0.56 s

OpenStudy (irishboy123):

apply in the vertical direction \(s = u_ot + \frac{1}{2} g t^2\) \(\implies s = \frac{1}{2} g t^2\) as \(u_0 = 0\) \(s = 0.9\) solve for \(t\).